A banquet hall has a 400 person capacity including both diners and servers. If one server is needed for every twelve diners, what is the maximum number of diners that can be served
Answer by richwmiller(17219)

400/13=30.x
30*13=390 total diners and servers
30*12=360 diners plus 30 servers
we still have room for 10 more
9 diners and 1 server
369 diners and 31 servers